t^5-256t=0
I took a t out and got: t(t4-256)=0
what do i do now?

Found 3 solutions by Alan3354, richwmiller, solver91311:
Answer by Alan3354(69443)

t^5-256t=0
-----------
t*(t^4 - 256) = 0
t = 0
------
(t^2 - 16)*(t^2 + 16) = 0
-----------------
t^2 - 16 = 0
t = +4
t = -4
-------
t^2 + 16 = 0
t = +4i
t = -4i

Answer by richwmiller(17219)

t^4=256
t^2*t^2=256
16*16=256
-16*-16=256
4,-4,4i,-4i are roots


Answer by solver91311(24713)



Well, one of your roots is obviously . Now you can then say that:



Let

Then you can say:



So



or



but since



which means that



or



which means that



Giving us the expected 5 roots for the initial quintic equation.

John